# Towards an exact electronic quantum many-body treatment of Kondo correlation in magnetic impurities

### v1

@@ Acknowledgements
Figure S5: Real-axis self-energies of Cr, Mn, and Fe impurities from all-orbital DMFT. (a) Real part of self-energies of $t_{2g}$ orbitals. (b) Real part of self-energies of $e_{g}$ orbitals. (c) Imaginary part of self-energies of $t_{2g}$ orbitals. (d) Imaginary part of self-energies of $e_{g}$ orbitals.
@@
We show real-axis self-energies of Cr, Mn, and Fe impurities in bulk Cu calculated by all-orbital DMFT in Fig. S5. When $\omega$ approaches the Fermi level, the imaginary part of the self-energies of the $3d$ orbitals approaches zero, while the real part of the self-energies of the $3d$ orbitals changes linearly with respect to the frequency, which agrees with the expected Fermi liquid behavior.
Supplementary Note 7 Summary of ground-state and spectral properties
The ground-state and spectral properties of magnetic impurities calculated by all-orbital DMFT are summarized in Table S4 and Table S5. In Table S5, we include the natural orbital occupancies of $t_{2g}$ and $e_{g}$ symmetry derived from the (36e, 52o) active-space DMRG density matrix. We choose to show the most fractional occupancies with $n_{\mathrm{nat}}>1$. We observe that the natural occupancies in Table S5 correlate well with the quasiparticle renormalization $Z$ in Table S4, i.e., more fractionally occupied orbitals are found to possess stronger quasiparticle renormalization.
Table S4: Hybridization function values at the Fermi level $\Delta(0)$, quasiparticle renormalization weights $Z$, and Kondo temperatures $T_{\mathrm{K}}$ of magnetic impurities in bulk Cu computed by all-orbital DMFT. All results are in the def2-SVP basis with Green’s function active space of (22e, 22o) ((22e, 27o) for Co).
@@
Table S5: Orbital occupancies $n$, natural orbital occupancies $n_{\mathrm{nat}}$, and spin moments $S$ of magnetic impurities in bulk Cu computed by all-orbital DMFT. All results are in the def2-SVP basis with ground-state active space of (36e, 52o) ((36e, 46o) for Ni).
@@
Supplementary Note 8 Spin fluctuation in Kondo impurities
@@
As shown in Fig. S6, Cr has the strongest intra- and inter-orbital spin correlation among all impurities. From Cr to Co, the intra-orbital spin correlation in $t_{2g}$ orbitals is largely suppressed, while the intra- and inter-orbital spin correlation within the $e_{g}$ orbitals stays strong or moderate, which is consistent with the orbital-dependent trend in quasiparticle renormalization. Fig. S6 also reveals that, in the Mn, Fe, and Co systems, the inter-orbital spin correlations have the following relation: $S_{i\neq j}(t_{2g}\text{-}t_{2g})<S_{i\neq j}(t_{2g}\text{-}e_{g})<S_{i\neq j}(e_{g}\text{-}e_{g})$.
Figure S6: Orbital-resolved spin-spin correlations for Cr, Mn, Fe, and Co impurities.
Supplementary Note 9 Effective $3d$-model calculations
To compare our all-orbital DMFT simulations against downfolded effective model calculations, we derived a multi-orbital Anderson impurity model with Hamiltonian
@@
The one-particle Hamiltonian in Eq. S12 was treated in a similar fashion as in the all-orbital DMFT calculations, where we adopted the Hartree-Fock effective Hamiltonian for five $3d$ orbitals so that the double-counting term can be exactly removed. The $3d$ block of the DFT hybridization function was discretized on the same non-uniform grid to obtain the bath parameters $\epsilon_{k}$ and $V_{ik}$, which resulted in 49 bath orbitals per $3d$ impurity orbital.
@@
To solve the embedding problem consisting of 5 impurity orbitals and 245 bath orbitals, we employed the same active-space DMRG solver. A Hartree-Fock calculation with fixed chemical potential at the DFT level was first performed, followed by a CISD calculation on the full embedding problem. A (40e, 40o) natural-orbital active space was derived by diagonalizing the CISD density matrix. We then conducted ground-state DMRG calculations with bond dimension $M=3500$ on the (40e, 40o) active space and further derived a (20e, 20o) DMRG natural-orbital active space. Finally, a dynamical DMRG calculation was done on the (20e, 20o) active space with bond dimension $M=1200$, and DMRG self-energies were used to estimate the quasiparticle renormalizations and Kondo temperatures in magnetic impurities. We summarize this result in Table S7 and the ground-state properties in Table S8. We find that the natural occupancies in downfolded model calculations are less fractional than the results in all-orbital DMFT calculations. In the meantime, the quasiparticle renormalization weights and the predicted Kondo temperatures are much higher in the model calculations compared to all-orbital DMFT.

### v3

@@ V $2\times 10^{-4}$ $7\times 10^{-4}$
Figure S5: Real-axis self-energies of Cr, Mn, and Fe impurities from all-orbital calculations. (a) Real part of the self-energy matrix elements diagonal in the orbital indices associated with the $t_{2g}$ orbitals. (b) Real part of the self-energy matrix elements diagonal in the orbital indices associated with the $e_{g}$ orbitals. (c) Imaginary part of self-energies of $t_{2g}$ orbitals. (d) Imaginary part of self-energies of $e_{g}$ orbitals.
@@
We show real-axis self-energies of Cr, Mn, and Fe impurities in bulk Cu calculated by all-orbital simulations in Fig. S5. When $\omega$ approaches the Fermi level, the imaginary part of the self-energies of the $3d$ orbitals approaches zero, while the real part of the self-energies of the $3d$ orbitals changes linearly with respect to the frequency, which agrees with the expected Fermi liquid behavior.
Figure S6: (a) $t_{2g}$ and (b) $e_{g}$ resolved DOS of Co impurity in bulk Cu from all-orbital calculations (“this work”), compared to DOS taken from Ref. [Surer2012a] (“Surer”).
Supplementary Note 7 Density of state of magnetic impurity orbitals
We present the local DOS of the $3d$ orbitals of the Co impurity in bulk Cu calculated by all-orbital simulations in Fig. S6, in addition to the DOS of Cr, Mn and Fe impurities in Figure 2. Compared to results from the 5-orbital model calculations in Ref. [Surer2012a], the all-orbital DOS exhibits qualitatively similar resonance and shoulder peaks near the Fermi level. Quantitatively, resonance peaks near the Fermi level from all-orbital calculations are narrower, especially for the $e_{g}$ orbitals. The narrower widths are consistent with the smaller quasiparticle renormalization weights $Z$ = 0.19 ($t_{2g}$) and 0.11 ($e_{g}$) from our all-orbital calculation compared to the $Z$ = 0.42 ($t_{2g}$) and 0.47 ($e_{g}$) from Ref. [Surer2012a].
Supplementary Note 8 Summary of ground-state and spectral properties
The ground-state and spectral properties of magnetic impurities calculated by all-orbital simulations are summarized in Table S6 and Table S7. In Table S7, we include the natural orbital occupancies of $t_{2g}$ and $e_{g}$ symmetry derived from the (36e, 52o) active-space DMRG density matrix. We choose to show the most fractional occupancies with $n_{\mathrm{nat}}>1$. We observe that the natural occupancies in Table S7 correlate well with the quasiparticle renormalization $Z$ in Table S6, i.e., more fractionally occupied orbitals are found to possess stronger quasiparticle renormalization.
Table S6: Hybridization function values at the Fermi level $\Delta(0)$, quasiparticle renormalization weights $Z$, and Kondo temperatures $T_{\mathrm{K}}$ of magnetic impurities in bulk Cu computed by all-orbital simulations. All results are in the def2-SVP basis with Green’s function active space of (22e, 22o) ((22e, 27o) for Co).
@@
Table S7: Orbital occupancies $n$, natural orbital occupancies $n_{\mathrm{nat}}$, and spin moments $S$ of magnetic impurities in bulk Cu computed by all-orbital simulations. All results are in the def2-SVP basis with ground-state active space of (36e, 52o) ((36e, 46o) for Ni).
@@
Supplementary Note 9 Spin and charge fluctuations in Kondo impurities
@@
As shown in Fig. S7a, Cr has the strongest intra- and inter-orbital spin correlation among all impurities. From Cr to Co, the intra-orbital spin correlation in $t_{2g}$ orbitals is largely suppressed, while the intra- and inter-orbital spin correlation within the $e_{g}$ orbitals stays strong or moderate, which is consistent with the orbital-dependent trend in quasiparticle renormalization. Fig. S7a also reveals that, in the Mn, Fe, and Co systems, the inter-orbital spin correlations have the following relation: $S_{i\neq j}(t_{2g}\text{-}t_{2g})<S_{i\neq j}(t_{2g}\text{-}e_{g})<S_{i\neq j}(e_{g}\text{-}e_{g})$. We also include the impurity charge fluctuation in Fig. S7b, which is calculated as $C_{ij}=\langle\hat{n}_{i}\hat{n}_{j}\rangle-\langle\hat{n}_{i}\rangle\langle\hat{n}_{j}\rangle$.
Figure S7: Orbital-resolved (a) spin and (b) charge correlations for Cr, Mn, Fe, and Co impurities.
Supplementary Note 10 Effective $3d$-model calculations
To compare our all-orbital simulations against downfolded effective model calculations, we derived a multi-orbital Anderson impurity model with Hamiltonian
@@
Table S11: Five-orbital model Hamiltonian results after fitting the chemical potential. Shown in the table are the changes in the chemical potential $\Delta\mu$, orbital occupancies $n$, quasiparticle renormalization weights $Z$, and Kondo temperature $T_{\mathrm{K}}$.
Impurity $\Delta\mu$ (eV) $n(t_{2g})$ $n(e_{g})$ $n(\mathrm{3d})$ $Z(t_{2g})$ $Z(e_{g})$ $T_{\mathrm{K}}(t_{2g})$ (K) $T_{\mathrm{K}}(e_{g})$ (K)
Ti 4.0 0.73 0.13 2.45 0.77 1.00 4171 5870
V 3.7 1.12 0.15 3.66 0.67 0.79 3183 4003
Cr 2.7 1.44 0.27 4.87 0.62 0.37 2649 1627
Mn 1.1 1.40 0.76 5.72 0.35 0.22 1411 864
Fe $-0.1$ 1.60 1.06 6.93 0.40 0.087 1547 315
Co $-1.9$ 1.56 1.62 7.93 0.47 0.78 1833 2719
Ni $-4.4$ 1.80 1.82 9.03 0.98 0.97 3963 3314
@@
The one-particle Hamiltonian in Eq. S13 was treated in a similar fashion as in the all-orbital calculations, where we adopted the Hartree-Fock effective Hamiltonian for five $3d$ orbitals so that the double-counting term can be exactly removed. The $3d$ block of the DFT hybridization function was discretized on the same non-uniform grid to obtain the bath parameters $\epsilon_{k}$ and $V_{ik}$, which resulted in 49 bath orbitals per $3d$ impurity orbital.
@@
To solve the embedding problem consisting of 5 impurity orbitals and 245 bath orbitals, we employed the same active-space DMRG solver. A Hartree-Fock calculation with fixed chemical potential at the DFT level was first performed, followed by a CISD calculation on the full embedding problem. A (40e, 40o) natural-orbital active space was derived by diagonalizing the CISD density matrix. We then conducted ground-state DMRG calculations with bond dimension $M=3500$ on the (40e, 40o) active space and further derived a (20e, 20o) DMRG natural-orbital active space. Finally, a dynamical DMRG calculation was done on the (20e, 20o) active space with bond dimension $M=1200$, and DMRG self-energies were used to estimate the quasiparticle renormalizations and Kondo temperatures in magnetic impurities.
We summarize the Green’s function results obtained using the same chemical potential as in the all-orbital calculations in Table S9 and the ground-state properties in Table S10. We find that the natural occupancies in the downfolded model calculations are less close to singly occupied than in the all-orbital calculations. In the meantime, the quasiparticle renormalization weights and the predicted Kondo temperatures are much higher in the model calculations compared to the all-orbital simulations. We also observe that the total $3d$ occupancies in the model differ from those in the all-orbital calculation by up to 1.27 electrons. Because the $3d$ occupancy can significantly impact the correlation strength and the Kondo temperature [jacob2015towards], for a fair comparison with the all-orbital calculations, we also performed the model calculation with chemical potentials $\mu$ fitted such that the total $3d$ occupancies match those in the all-orbital calculations in Table S7. The model results with the fitted $\mu$ are shown in Table S11. The Kondo temperatures calculated using the new quasiparticle renormalization weight decrease slightly for a few elements (e.g., Fe and Co) compared to without the chemical potential fitting, but there is no consistent improvement in the absolute $T_{\mathrm{K}}$ or its trends. With or without fitting the chemical potential in the low-energy model calculations, our all-orbital results are an order of magnitude better.
Supplementary Note 11 Specific heat calculation on the Anderson impurity model
We benchmark the accuracy of a specific heat calculation on the symmetric single-impurity Anderson model (SIAM) (see Supplementary Note 2 for an introduction) within the numerical framework of this work. In the symmetric SIAM, for large $\frac{U}{\Delta_{0}}$, $T_{\mathrm{K}}$ can be calculated analytically as a function of the impurity on-site Coulomb interaction $U$ and the hybridization strength $\Delta_{0}$, $T_{\mathrm{K}}=\sqrt{\frac{U\Delta_{0}}{2}}\exp\left(-\frac{\pi U}{8\Delta_{0}}+\frac{\pi\Delta_{0}}{2U}\right)$  [Hewson1993a]. Additionally, the specific heat has been solved exactly through Bethe Ansatz and computed numerically [wiegmann1983exact, okiji1983thermodynamic]. This allows us to analyze the errors due to the bath discretization and active space truncation on the peak temperature of the specific heat and also verify the nature of the excitations that give rise to the Kondo singlet breaking (see Supplementary Note 12).
To compute the finite temperature properties, we solve for the excited states of the SIAM in the canonical ensemble using the ground-state DMRG solver with SU(2) symmetry and a level-shifted Hamiltonian for the $n^{\mathrm{th}}$ eigenstate
$\hat{H}^{\prime}_{n}=\hat{H}+\sum_{i=1}^{n-1}c_{i}|\Psi_{i}\rangle\langle\Psi_{i}|$ (S20)
where $\hat{H}$ is the original DMRG Hamiltonian, $c_{i}$ is the energy shift multiplier, and $|\Psi_{i}\rangle$ is the $i^{\mathrm{th}}$ converged eigenstate to be projected out [Zhai2023a, wouters2014chemps2, fishman2022itensor]. We used $c_{i}=5D$ for all $i$ and a DMRG bond dimension of $M=1200$. With this bond dimension, DMRG converged to a small discarded weight of less than $5\times 10^{-9}$.
Figure S8: Benchmark of finite temperature results from the DMRG solver on the SIAM with $U=0.012D$ and $\Delta_{0}=0.001D$, using the logarithmic discretization with base $b=4$ and the procedure to extract impurity specific heat as outlined in Ref. [merker2012numerical]. Different components of the energy and specific heat are defined the same as in Fig. 4 of Ref. [merker2012numerical]. The exact $T_{\mathrm{K}}$ (vertical black dotted line) and specific heat peak temperature from numerical renormalization group results in Ref. [merker2012numerical] (vertical black dashed line, “Merker”) are included as references.
We benchmarked on the SIAM with the same parameters as in Ref. [merker2012numerical], $U=0.012D$ and a flat-band hybridization $\mathrm{Im}\Delta(\omega+i0^{+})=-\Delta_{0},\leavevmode\nobreak\ |\omega|<D$ where $\Delta_{0}=0.001D$. For convenience, we set $D=1$. We used the same logarithmic discretization, consistent with the rest of the manuscript, but with 19 orbitals and logarithmic intervals defined using a larger log base $b=4$ as used in Ref. [merker2012numerical] to achieve a small spacing near the Fermi level. Using the procedure to extract the impurity specific heat contribution as outlined in Ref. [merker2012numerical], we obtained a peak in the impurity specific heat at 0.22 $T_{\mathrm{K}}$ (Fig. S8), in good agreement with the peak in Ref. [merker2012numerical] at around 0.28 $T_{\mathrm{K}}$.
In Fig. S9 we show the estimated specific heat peak temperatures $T_{C_{\mathrm{V}}}$ as a function of the (DMRG natural orbital) active space size. We see that the position of the peak varies by about a factor of 3 as a function of the active space size (and thus the estimated Kondo temperature also varies by about a factor of 3). This is different from the behaviour of the estimated Kondo temperatures from the Hewson formula, as shown in Supplementary Note 2, which are already reasonably converged with an active space size of 8 orbitals. Since the location of the $C_{\mathrm{V}}$ peak depends on different excitations (in particular the neutral excitations) than those that enter into the self-energy and $Z$, it is unsurprising that there are different finite size effects. These numerical results support the use of $T_{\mathrm{K}}$ from the Hewson formula as our primary estimate in the main text.
Figure S9: Specific heat peak temperature ($T_{C_{\mathrm{V}}}$) of SIAM with $U=0.012D$ and $\Delta_{0}=0.001D$, using active spaces of different numbers of orbitals taken out of the full space of 20 orbitals, the logarithmic discretization with base $b=4$ and the procedure to extract the impurity specific heat as outlined in Ref. [merker2012numerical]. The exact $T_{\mathrm{K}}$ (black dotted line) and specific heat peak temperature from numerical renormalization group results in Ref. [merker2012numerical] (black dashed line) are shown as references.
From the excited states, we can also examine the nature of the excitations that give rise to the peak in the heat capacity. To verify the spin-flip nature of the lowest energy excitations, we compute the impurity spin-flip weight $w_{ii}^{2}=|\langle\Psi_{S=0}|c^{\dagger}_{i\beta}c_{i\alpha}|\Psi_{S=1}\rangle|^{2}$ where $i$ indexes the impurity orbital (here, the $S_{z}=1$ state of the $|\Psi_{S=1}\rangle$ multiplet is used as an example). This quantity satisfies $w_{ii}^{2}\leq 1$ and is equal to 1/2 for a pure localized impurity spin flip. We show the weights for the lowest excited states, with their corresponding energies, in Table S12.
Table S12: Impurity spin-flip weight $w^{2}_{ii}$, its normalized value $w^{2}_{ii}/\sum_{pq}w^{2}_{pq}$, and the excitation energies $E/D$ of the lowest singlet-triplet excitations that give rise to the specific heat peak, calculated in active spaces of different number of orbitals ($N_{\mathrm{orb}}$).
$N_{\mathrm{orb}}$ threshold (occupied) threshold (virtual) $w^{2}_{ii}$ $w^{2}_{ii}/\sum_{pq}w^{2}_{pq}$ $E/D$
20 0 0 0.071 0.123 $1.2\times 10^{-5}$
14 $1\times 10^{-7}$ $1\times 10^{-7}$ 0.073 0.105 $1.2\times 10^{-5}$
10 $3\times 10^{-5}$ $3\times 10^{-5}$ 0.087 0.138 $1.5\times 10^{-5}$
8 $2\times 10^{-4}$ $2\times 10^{-4}$ 0.101 0.169 $2.0\times 10^{-5}$
6 $2\times 10^{-3}$ $2\times 10^{-3}$ 0.121 0.210 $3.4\times 10^{-5}$
Supplementary Note 12 Ab initio impurity specific heat
We calculated the finite-temperature and excited state properties in the all-electron simulations within different active spaces, using the procedure outlined in Supplementary Note 11. We used a grand canonical formalism and extracted the canonical $C_{\mathrm{V}}$ using the relations in Ref. [stroker2021classical]. The chemical potential was chosen to ensure that the average electron number of the impurity and bath at $T_{\mathrm{K}}$ (estimated from Hewson’s formula Eq. 3) matched the electron number of the ground state. The excited states were obtained in a DMRG active space of up to (26e, 36o) size with maximum DMRG bond dimension $M=4000$, converging to a discarded weight of less than $5\times 10^{-5}$. In this way, 449, 129, 83, 44, 42, 38, and 80 excited states in total were calculated for the Ti, V, Cr, Mn, Fe, Co, and Ni impurities, respectively.
In the SIAM, the impurity contribution to the specific heat was obtained by partitioning the energy between the impurity and the bath, as described in Ref. [merker2012numerical]. However, in the case of the ab initio model, the partitioning is ambiguous. This is in part because energy partitioning is always non-unique (as the only constraint is that the parts sum up to the same whole), but also because as we converge to the basis limit of the all-electron basis, the additional impurity orbitals necessarily extend beyond the impurity atom into the bulk (see discussion in main text). We found that this either made the estimated $C_{\mathrm{V}}$ sensitive to various choices, or gave unphysical results (e.g., the impurity specific heat capacity could be negative).
Consequently, to produce a (rough) estimate of $T_{\mathrm{K}}$, similar to the analysis of the excitation character of the excited states in Supplementary Note 11, we quantified the degree of “impurity” character of each excited state and only included those above a given threshold when computing the impurity specific heat. Concretely, for the $S_{z}=1$ component of a triplet excited state, we computed
$w_{ij}=\langle\Psi_{0}|c^{\dagger}_{i\alpha}c_{j\beta}|\Psi_{1}\rangle$ (S21)
where $c^{\dagger}_{i\alpha},c_{j\beta}$ are creation and annihilation operators on orbitals that are localized on the impurity or on the bath, respectively (these orbitals were obtained by a rotation of the active space). Then, the impurity character is defined as
$\lambda=\frac{\sum_{i\notin\mathrm{bath}\ \mathrm{or}j\notin\mathrm{bath}}w_{ij}^{2}}{\sum_{ij}w_{ij}^{2}}.$ (S22)
A similar analysis was carried out for the quintet excited states. Using a threshold of $\lambda>0.12$, we show the estimated $T_{C_{\mathrm{V}}}$ from the impurity specific heat peak in Fig. S10. In the SIAM, the peak of the specific heat is always below $T_{\mathrm{K}}$. We see that this is true in our calculations for Ti and V, but not for the other impurity elements, suggesting that $T_{\mathrm{C_{\mathrm{V}}}}$ is overestimated in the other cases. Given the strong finite size effects observed in specific heat capacity calculations in general (as studied in the SIAM in Supplementary Note 11), we consider the estimation of $T_{\mathrm{K}}$ from the Hewson formula to be better converged. However, this rough treatment indicates that similar trends can be reproduced through the heat capacity (see Fig. S11).
Figure S10: Impurity specific heat per mole of magnetic impurities in bulk Cu from ab initio all-orbital calculations. The Kondo temperatures predicted using Herson’s formula with the quasiparticle renormalization weights from the Green’s function calculations (“$T_{\mathrm{K}}$ from GF”) are shown for comparison.
Figure S11: The temperatures corresponding to the peak in the impurity specific heat compared to the Kondo temperatures predicted by Hewson’s formula from ab initio all-orbital calculations and those measured in experiments.
Supplementary Note 13 Ab initio and model excited state analysis
To illustrate the difference in the excited states in the ab initio all electron treatment and in the model calculations, we performed excited state model calculations for Mn in the (20e, 20o) active space with the same chemical potential as in the all-orbital calculations, and compared those to the excited states seen in the ab initio calculations. As shown in Fig. S12, in both the model and ab initio all-orbital calculations, we observed multiple near-degenerate low-energy triplets that involve partial transfer of an electron from $t_{\mathrm{2g}}$ to $e_{g}$ and to the bath orbitals, accompanied by a spin-flip (orange), as well as one low-energy triplet that corresponds to a simultaneous on-site spin-flip of all fractionally occupied $d$ orbitals (purple). The latter corresponds to the low-energy excitation of a large spin Kondo model. However, in the model calculations, the charge transfer triplets are higher in energy than the Kondo model triplet, which is opposite to the all-orbital calculations where the charge-transfer triplets are the lowest triplets. As discussed in the main text, correctly capturing the charge-transfer triplet excitations appears to be critical to reproducing the quantitative Kondo trends, which highlights again the power of the all-orbital calculations.
Figure S12: The excitation energies of triplets associated with on-site spin-flip (purple) from the ground state and of triplets associated with charge transfer between different orbitals and spin-flip (orange) from the ground state for the Mn impurity. The impurity-impurity and impurity-bath blocks of spin-flip transition density matrices $|\langle\Psi_{S=0}|c^{\dagger}_{i\downarrow}c_{j\uparrow}|\Psi_{S=1,S_{z}=1}\rangle|$ from the ground state (vertical axis) to the triplet(s) (horizontal axis) from model calculations and ab initio all-orbital calculations are shown on the left and right, respectively, to illustrate the on-site spin-flip (purple) and the charge transfer and spin-flip (orange) process. The bath orbitals are ordered based on their bath energies $\omega$. The transition density matrices of charge-transfer triplets are summed over all near-degenerate states.
Figure S13: Orbital resolved spin-flip transition density matrix $|\langle\Psi_{S=0}|c^{\dagger}_{i\downarrow}c_{j\uparrow}|\Psi_{S=1,S_{z}=1}\rangle|$ between the ground state and the lowest triplet excited state that is dominated by intra-impurity orbital excitations, across the series of impurities from Ti to Co.
Fig. S13 further shows the ab initio transition density matrix between the ground-state and the low-energy triplet state which is dominated by a pure intra-impurity spin-flip for all the elements, corresponding to the excitation in the multi-orbital Kondo model. For Ti and V, the singlet-triplet excitation is dominated by the $t_{2g}$ intraorbital excitations, while for Fe and Co it is dominated by the $e_{g}$ intraorbital excitations. The excitations in Cr and Mn have significant contributions from both $t_{2g}$ and $e_{g}$ orbitals.
Supplementary Note 14 Computational time
Table S13: Computational time of various tasks: ground state DMRG calculations in the largest active space (36e, 52o); Green’s function calculations with DDMRG at 7 frequency points near the Fermi level in active spaces of around (22e, 22o) ((22e, 27o) for Co); and specific heat ($C_{V}$) calculations where we calculated ground and excited eigenstates in spin manifolds of up to $S=8$ in the canonical ensemble in active spaces of around (28e, 28o) ((30e, 29o) for Ti, (28e, 30o) for V, (26e, 36o) for Co). The ground state and Green’s function calculations were performed on 2–8 nodes with 24 cores per node, while the specific heat calculations in each spin manifold were performed on 1–8 nodes with 64 cores per node.
Ground state (36e, 52o) Green’s function (22e, 22o) $C_{V}$ (28e, 28o)
Impurity wall time (h) core hours wall time (h) core hours wall time (h) core hours
Ti 27 2556 60 5776 1019 119784
V 55 7984 94 18057 733 95274
Cr 64 9215 99 18959 645 50843
Mn 100 14395 105 20088 489 49726
Fe 82 11853 135 25848 237 15174
Co 78 11225 298 71534 171 21884
Ni 49 9407 66 12753 50 3179
Supplementary Note 15 Kondo temperature in multi-orbital impurities
The exponential decrease in the Kondo temperature with the spin $S$ of the impurity [Daybell1968] has previously been discussed using a mechanism involving Hund’s coupling [Schrieffer1967]. We briefly summarize the literature arguments here.
As a simple example, consider an Anderson model with $l=2$ conduction electrons scattering off an impurity with a half-filled $d$-shell,
$H=\sum_{\bm{k}m\sigma}\epsilon_{\bm{k}}n_{\bm{k}m\sigma}+\sum_{\bm{k}m\sigma}V_{\bm{k}m}\left(c^{\dagger}_{\bm{k}\sigma}f_{m\sigma}+f^{\dagger}_{m\sigma}c_{\bm{k}\sigma}\right)+H_{\mathrm{imp}}\text{,}$ (S23)
where $H_{\mathrm{imp}}$ includes the interaction on impurity sites and $m,\sigma$ are the labels of the atomic orbital and spin separately.
In the presence of strong on-site electron repulsion and Hund’s coupling, we can focus solely on the spin degree of freedom of the impurity. The spin configuration of the impurity is constrained to $S=5/2$ Using the Schrieffer-Wolff transformation [Schrieffer1966, Schrieffer1967], the effective Hamiltonian is reduced to the following form:
$H=\sum_{\bm{k}m\sigma}\epsilon_{\bm{k}}n_{\bm{k}m\sigma}+\sum_{\bm{k}\bm{k}^{\prime}m}J_{m\bm{k}\bm{k}^{\prime}}\bm{S}\cdot c^{\dagger}_{\bm{k}^{\prime}m}\frac{\bm{\sigma}}{2}c_{\bm{k}m}\text{.}$ (S24)
The orbital angular momentum of the scattering conduction electrons is conserved. Because the impurity is half-filled, conduction electrons cannot exchange orbital angular momentum with the impurity. The exchange interaction $J_{m\bm{k}\bm{k}^{\prime}}$ is second order in the hybridization term:
$J_{m\bm{k}\bm{k}^{\prime}}=\frac{V_{\bm{k}m}V_{\bm{k}^{\prime}m}}{2S}(\frac{1}{E_{+}}+\frac{1}{E_{-}})\text{,}$ (S25)
where $E_{+}$ and $E_{-}$ represent the additional energy of the virtual process involving the addition or removal of an electron. The factor of $2S$ in the denominator arises from the representation of the $S=5/2$ spin operator in terms of five $S=1/2$ impurity electron operators.
For an impurity without a half-filled $d$-shell, the effective Hamiltonian includes both orbital and spin exchange. However, if we focus only on the spin degree of freedom and assume that all singly occupied states are spin-up, with the remaining states either doubly occupied or empty, the spin dependence of $J_{m\bm{k}\bm{k}^{\prime}}$ can be generalized, where spin corresponds to the spin of the singly occupied orbitals [Schrieffer1967]. Although simplistic, this formula provides a rationalization for the trend of the Kondo temperature $T_{\mathrm{K}}$ of transition metal impurities within host alloys [Daybell1968, Nevidomskyy2009] with
$T_{\mathrm{K}}\sim\exp(-1/J^{\mathrm{eff}}\rho)\text{.}$ (S26)
Here, $J^{\mathrm{eff}}$ is inversely proportional to the spin $S$ as in Eq. S25, and consequently, the Kondo temperature is predicted to exhibit an exponential dependence on $S$.
Another way to estimate the Kondo temperature is through the bandwidth of the Kondo resonance (Abrikosov-Suhl resonance), $\tilde{\Delta}$ , at the Fermi level, with
$T_{k}=\frac{\pi}{4}\tilde{\Delta}\text{.}$ (S27)
If there is no quantum phase transition between the non-interacting limit and the finite-interaction case, the low-energy behavior of the interacting system can be described by a local Fermi liquid theory [Nozieres1974]. The bandwidth $\tilde{\Delta}$ in the interacting case is related to the non-interacting hybridization width $\Delta$ through the quasiparticle renormalization weight $Z$ :
$\tilde{\Delta}=Z\Delta\text{.}$ (S28)
The entire effect of the interaction is encapsulated in $Z$. If the renormalization weight $Z$ can be estimated non-perturbatively, this formula can be applied to address the finite interaction case, which is the basis of the Hewson estimate of the Kondo temperature used in the main text.
[…]
[title=References in Supplementary Materials, notkeyword=appendix]